OpenStudy (anonymous):

BCD is a parallelogram. If angle B is equal to (2x)° and angle A is equal to (-x + 120)°, find x. I got x=60. Is it correct?

OpenStudy (danjs):

the angles are equal it sounds like, opposite angles in a parallelogram

OpenStudy (danjs):

2x = -x+120

OpenStudy (danjs):

so you got it !

OpenStudy (anonymous):

ABCD is a parallelogram. Sorry the A was missing.

OpenStudy (anonymous):

wouldn't 2x = -x+120 give 40?

OpenStudy (danjs):

one sec---

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\[2X+(-X+120)=180\]

OpenStudy (danjs):

they gonna add up to 180

OpenStudy (anonymous):

So 2X+(−X+120)=180 = 2x-x+120 = 180 x+120=180 x=60??

OpenStudy (danjs):

yes

OpenStudy (anonymous):

Thanks :D

OpenStudy (danjs):

np, ask whenever, @DanJS